If a = b^2/b-a then find the value of  a^3 + b^3
Please help me out

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
As 
a = b² /(b-a)
Hence
ab - a² = b²
a² + b² -ab = 0
Now
a³ + b³ = (a+b) (a² + b² -ab) = (a+b)* 0 = 0
Hope it helps
 
Regards
Arun (askIITians forum expert)